# Non-local, diamagnetic electromagnetic effects in long MITLs

### v1

@@ V Magnetic insulation model
The issue at hand is the value of the SCL current crossing the AK gap as a function of magnetic field magnitude. A theoretical value of the critical magnetic field (also called the Hull field), at which complete insulation occurs, was derived by Lovelace and Ott lovelace_theory_1974 . The expression they derived depends on the value of the potential (difference), $V_{0}$, [cf. Eqs. (12) and (13)] and is given by
@@
The expression (19) was experimentally tested by Orzechowski and Bekefi orzechowski_current_1976 who found that magnetic insulation occurred mostly in agreement with Eq. (19) but SCL current did not sharply drop to zero at the Hull magnetic field value; instead, it had a “spillover.” They constructed a curve by normalizing the measured current to $V_{0}^{3/2}$ (from the Child-Langmuir law) and the magnetic field to the critical magnetic field, (19). We follow Ref. vandevender_requirements_2015 and normalize the loss current to the Child-Langmuir current. Since the expression for the Child-Langmuir current folds in the particular geometry, one can think of such a curve as universal. We refer to this curve as the Hull curve.
@@
As CEM is concerned in regards to predicting SCL current losses in MITLs, we note that major codes, such as BERTHA and SCREAMER, use the curve plotted in Fig. 7 of Ref. vandevender_requirements_2015 , labeled “Original SCREAMER.” Reference vandevender_requirements_2015 also proposes a revised version of the curve, however, such revision does not appear to have been widely implemented spielman_private_2024 . We refer to the curve “Original SCREAMER” from Ref. vandevender_requirements_2015 more generally as the “Circuit Element fit.” This curve is shown as a dashed line [^3] Our fit of that curve is based on the image from Ref. vandevender_requirements_2015 . in Fig. 7. Reference vandevender_requirements_2015 additionally states that the Hull curve has been constructed from 3D PIC simulations. We attempted to replicate the curve using 2D PIC simulations with CHICAGO and found that our results did not agree with the Circuit Element fit curve, as indicated by the black triangles in Fig. 7.
@@
The source of disagreement is worth discussing. One can see that our simulation data stops at about $B/B_{\rm crit}=0.91$. The reason behind this is that as we approached the Hull magnetic field, the electron layer transitioned from stable and steady-state to unstable. In an unstable layer, the current collected by the anode as well as the electric field acquire oscillatory behavior (the magnetic field stays nearly constant since it is externally imposed). In such a system, averaged values are difficult to obtain due to the randomness in the oscillations. Even if averaged values were possible to reliably obtain, they would not be representative of a steady-state system, upon which assumption both the Child-Langmuir law and Hull magnetic field calculations are based. However, knowing an accurate value of $V_{0}$ is crucially important since both the Child-Langmuir current and the critical magnetic field strongly depend on it, as seen from Eqs.(12) and (19). These two quantities are being used to normalize the measured current and magnetic field, and because of that, any uncertainty in $V_{0}$ translates into uncertainty in the normalized quantities on the Hull curve. Our experience has shown that near cutoff, small uncertainty in $V_{0}$ can easily place a point on either side of the theoretical Hull field.
@@
Apart from the uncertainty associated with the value of $V_{0}$, our PIC simulations also show the “spillover” effect and loss of a sharp insulating boundary. These observations have led us to revisit the interpretation of the experimental results in Ref. orzechowski_current_1976 . In that reference, it was hypothesized that some process, not characterized by the authors, produced “hot tail” electrons that contributed to smearing of the sharp theoretical boundary. The authors detected radiation when the magnetic field strength exceeded the Hull field and no such radiation for lower magnetic fields, which they also correlated with the onset of an instability in the system. To our knowledge, the precise mechanism for such “hot electrons” has not been identified in the literature to date. Here we suggest another possible instability leading to the loss of a sharp insulating boundary, in agreement with our PIC simulations. Namely, we propose that the electron layer loses stability due to a velocity shear type instability in crossed field devices, such as diocotron, magnetron, etc. In such an unstable layer, our observations show the formation of large-scale vortical structures that move in a stochastic way throughout the system (largely because of $\mathbf{E}\times\mathbf{B}$ drift), leading to stochastic oscillatory behavior in all quantities within the AK gap, including in the conducting current. Such mechanism results in a loss current beyond the Hull field but does not necessarily produce a population of hot electrons. Ultimately, to be able to unambiguously identify the precise instability, a further study beyond the scope of this work is necessary.
@@
The stability properties of a parapotential magnetized electron layer have been discussed in the past buneman_stability_1966 ; davidson_equilibrium_1991 ; davidson_physics_2001 . While these properties depend on the layer’s spatial profile, of equal importance is whether and how this profile may in practice (experiment) be achieved. Many factors make this problem very complex, however, for relevant to $Z$ parameters, even the simplest constant profile electron layer is never strictly stable, something recently emphasized in Ref. evstatiev_efficient_2023 . An attempt to create a stable electron layer by the authors of Ref. bergeron_beam_1979 was also not successful. We believe that in order to properly understand the loss of a sharp insulating boundary, a theory of current transport across an AK gap in unsteady flows must be developed.
@@
with positive $a$, $b$, and $\alpha$. Our Analytic fit Hull curve is obtained for values $a=0.065$, $b\in[1.03,1.04]$, and $\alpha=1$; the Circuit Element fit is obtained for $a=0.78$, $b=1.4$, and $\alpha=4$. The notable feature here is that our analytic fit has a much smaller range outside the critical magnetic field, by only about 33334444%, compared to about $40$% for the Circuit Element fit.
Figure 8: Comparison of current losses in a 555\,5m long MITL with $I_{\rm peak}=20\,$MA, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm. The Circuit Element fit overpredicts EMPIRE 2D PIC result by about a factor of $3$ while the Analytic fit is in excellent agreement with it.
Since our PIC simulations were not successful in constructing the entire Hull curve, we have fitted the parameters in Eq. (20) from PIC simulations of current loss in MITLs of various geometrical dimensions. In particular, long MITLs have an amplifying effect on small variations in parameter values. Fig. 8 shows a simulation of current loss in a 555\,5m long MITL ($I_{\rm peak}=20\,$MA, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm), which illustrates the difference in the predictions using the two different fits from Fig. 7. Clearly, the Circuit Element fit greatly overpredicts current losses, by a factor of about $3$. (We have confirmed this result also by a direct simulation with BERTHA, not shown here.) Conversely, the details of the Hull curve in shorter MITLs in the length range $\lesssim 2\,$m do not manifest significant discrepancy with PIC simulations and both fits produce acceptable predictions.
One last remark in this section is that a “spillover” Hull curve acts as a low pass filter. Looking back at Fig. 5, we notice that the 1D model contains a certain amount of small scale temporal oscillations. Such oscillations are not PIC noise since the 1D model is a (cold) fluid model. Comparing now the simulation with the Circuit Element fit in Fig. 8, we see a noticeably smoother curve. This is not to imply that removing the small scale oscillations is unphysical, rather, that it is possible that certain physical oscillations may unintentionally be filtered out.
@@ VI Temperature rise due to electron impact
We use the NIST stopping power tables for electrons in metals, $dK/dx$ nist_estar ; or for certain metals, we prefer to use an analytic expression that gives the extra information in the low energy range nguyen-truong_modified_2015 . First, the electron flux to the anode is directly related to the electron current:
@@
where $M_{\rm mol}$ is the molar mass of the metal, $\rho$ is the mass density, $C_{V}$ is the heat capacitance at constant volume, and $\theta_{i}$ is the electron angle of incidence measured from the normal to the surface (see also Ref. bennett_current_2019 ). The total temperature increase is the sum over all time steps, $N_{\rm ins}=t_{\rm ins}/\Delta t$, until complete magnetic insulation is achieved,
@@
The only extra information necessary to calculate the temperature rise is the angle of incidence, which can be calculated as follows. Using the steady-state relations for conservation of the $z$-component of canonical momentum and the energy balance in Cartesian geometry lovelace_theory_1974 ; bergeron_relativistic_1975 , we have
@@
Upon using Eq. (25), it is easy to recognize that setting the radical expression in Eq. (28) to zero, i.e., $\gamma^{2}-1-\left(eB_{y}d/mc\right)^{2}=0$, amounts to precisely the Hull condition [^5] The Hull condition can be derived from the same equations by setting $v_{x}=0$ at $x=d$, i.e., the condition that electrons do not reach the anode. (19). This means that the angle of incidence (28) is undefined at insulation (by Lovelace and Ott) as well as for larger values of the magnetic field, for which the expression under the square root becomes negative. In other words, using a Hull curve with a “spillover” is incompatible with the so outlined method of calculating the temperature. The simplest approach to resolving this incompatibility is to neglect temperature contributions at locations where $B_{y}\geq B_{\rm crit}$. This simple solution is acceptable for two reasons. First, we have seen from the Hull curve in Fig. 7 that the electron flux, which is proportional to the current, significantly drops near $B_{\rm crit}$, and so does its contribution to the temperature rise. And second, the neglected range of magnetic fields is only a small fraction of 33334444% of the total range for our Analytic fit. Our practical experience has confirmed that temperature doesn’t noticeably increase at locations near magnetic insulation.
@@
As a verification of this approach, Fig. 9 shows the temperature rise calculated from data from our 1D model and from CHICAGO 2D PIC, for aluminum metal initially at 300300300\,300K. The simulation setup is the same as in Fig. 4. To avoid small differences due to algorithmic implementation details, we use CHICAGO’s output quantities of current density and kinetic energy near the anode (quantities from particles information accumulated to the grid) to perform the calculations leading to Eq. (23). We notice that the overall agreement between the two models is excellent. This implies also that not only do both components of the current density agree very well (we have not shown $z$-components form either model) but so do the kinetic energies. The slight disagreement in the range of 20202020303030\,30cm at $t=9\,$ns (top panel) is due to the slight temporal delay of the PIC pulse compared to the 1D model, see Fig. 4 (bottom panel); however, just 111\,1ns later, at 101010\,10ns (bottom panel) no such disagreement is seen (the current loss pulses have caught up to one another). Since the temperature calculation is cumulative, such temporal offsets would not affect the final temperature increase.
@@
The obvious question concerns the large temperature oscillations in the PIC simulation in the first $\sim 10\,$cm of the MITL. These oscillations correspond to the current (and flux) oscillations in Fig. 4, which we believe to be a numerical artifact of the SCL emission algorithm in PIC, as discussed in Ref. evstatiev_efficient_2023 .
@@
Figure 9: Comparison of temperature increase at $t=9\,$ns (top) and 101010\,10ns (bottom) between the 1D model and CHICAGO 2D PIC simulations, for $I_{\rm peak}=20\,$MA, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm, and $L=1\,$m.
@@ VII Conclusions
In this work we have discussed MITL losses in a strongly time-dependent setting. A major effect that helps dramatically reduce current losses in long MITLs is that of the diamagnetic, non-local cumulative electromagnetic effect of SCL currents. In short MITLs these effects diminish and in the limit vanish.
A new 1D electromagnetic model is developed that performs six to seven orders of magnitude more efficiently than 2D PIC computations. The model has been verified against 2D PIC simulations of parallel plate Cartesian and straight coaxial MITL with azumuthal symmetry. It allows for efficient navigation of the large MITL design parameter space. A number of scaling studies have been performed, helping to understand and minimize MITL losses. A temperature diagnostic within the 1D model is easy to implement, based on the available information from the fields and currents. Agreement with the temperature calculations from PIC data further verifies that the 1D model accurately predicts kinetic energies as well.
A revised magnetic insulation model is proposed in the form a Hull curve, partly informed by PIC simulations and partly by fitting PIC simulation data. It is found that close to magnetic insulation, the electron layer between the anode and cathode becomes unstable and prevents the computation of the complete Hull curve. Major circuit element codes using a previously computed Hull curve tend to overpredict current losses in long MITLs but implementation of our revised Hull curve should not present difficulty.
It is proposed that previous experimental observations by Orzechowski and Bekefi of non-zero loss current beyond the theoretical critical magnetic field by Lovelace and Ott is due to the electron layer losing stability, which also fits within the physics picture based on PIC simulations. This hypothesis still needs experimental confirmation.

### v2

@@ V Magnetic insulation model
The issue at hand is the value of the SCL current crossing the AK gap as a function of magnetic field magnitude. A theoretical value of the critical magnetic field (also called the Hull field or Hull cutoff [7]), at which complete insulation occurs, was derived by Lovelace and Ott [8]. (An earlier calculation, based on a single particle motion in external fields, was given by Walker [44].) The expression they obtained depends on the value of the potential (difference), $V_{0}$, [cf. Eqs. (12) and (15)] and is given by
@@
This dependence was experimentally tested by Orzechowski and Bekefi [45] who found that magnetic insulation occurred mostly in agreement with Eq. (20) but that the SCL current did not sharply drop to zero at the Hull magnetic field value; instead, it had a “spillover” with non-zero current reaching the anode past the critical cutoff. Reference [45] presents curves of SCL current vs. magnetic field in normalized units, whereby their measured current was normalized to either $V_{0}^{3/2}$ (termed perveance) or to the Child-Langmuir current; they normalized the measured magnetic field to the theoretical critical magnetic field(20). In the following, we normalize our SCL currents to the Child-Langmuir current and magnetic fields to $B_{\rm crit}$. We refer to the curve of normalized emitted SCL current vs. normalized magnetic field as the Hull curve. Since the expression for the Child-Langmuir current folds in the particular geometry, one can think of such a curve as universal.
@@
As CEM is concerned in regards to predicting SCL current losses in MITLs, we note that major codes, such as BERTHA and SCREAMER, use the curve plotted in Fig. 7 of Ref. [46], labeled “Original SCREAMER.” Reference [46] also proposes a revised version of the curve, however, such revision does not appear to have been widely implemented [47]. We refer to the curve “Original SCREAMER” from Ref. [46] more generally as the “Circuit Element fit.” This curve is shown as a dashed line [^3] Our fit of that curve is based on the image from Ref. [46]. in Fig. 7. Reference [46] additionally states that the Hull curve has been constructed from 3D PIC simulations. We attempted to replicate the curve using 2D PIC simulations with CHICAGO and found that our results did not agree with the Circuit Element fit curve, as indicated by the black triangles in Fig. 7.
@@
One aspect of the disagreement is worth discussing. One can see that our simulation data stops at about $B/B_{\rm crit}=0.91$. The reason behind this is that as we approached the Hull magnetic field, the electron layer transitioned from stable and steady-state to unstable. In an unstable layer, the current collected by the anode as well as the electric field acquire oscillatory behavior. (The magnetic field is much more constant since it is externally imposed and typically much larger than the (self-)magnetic field of the SCL currents. There is no contradiction with the results in the previous sections since here the electromagnetic effects were negligible due to the system being in either a steady state or having slow fluctuations when unstable. In addition, the system length in these simulations was quite short, about 101010\,10cm, which also necessitated using an artificially low threshold value of 380380380\,380kV/m.) In such an unstable system, averaged values are difficult to obtain due to the randomness in the oscillations. Even if averaged values were possible to reliably obtain, they would not be representative of a steady-state system, upon which assumption both the Child-Langmuir law and Hull magnetic field calculations are based. However, knowing an accurate value of $V_{0}$ is crucially important since both the Child-Langmuir current and the critical magnetic field strongly depend on it, as seen from Eqs.(12) and (20). These two quantities are being used to normalize the measured current and magnetic field, and because of that, any uncertainty in $V_{0}$ translates into uncertainty in the normalized quantities on the Hull curve. Our experience has shown that near cutoff, even a small uncertainty in $V_{0}$ can easily place a point on either side of the theoretical Hull field.
@@
Apart from the uncertainty associated with the value of $V_{0}$, our PIC simulations  [^4] Our simulations with CHICAGO were electromagnetic and a wave was launched into the simulation domain using a simple input circuit. The value of the electric potential $V_{0}$ entering Eqs. (12), (15), (20) was calculated using relation (17) with $\widetilde{E}_{x}(x=d;z,t)$ provided by CHICAGO. confirm the “spillover” effect and the loss of a sharp insulating boundary. These observations have led us to revisit the interpretation of the experimental results in Ref. [45]. In that reference, it was hypothesized that some process, not characterized by the authors, produced “hot tail” electrons that contributed to smearing of the sharp theoretical boundary. The authors detected radiation when the magnetic field strength exceeded the Hull field and no such radiation for lower magnetic fields, which they also correlated with the onset of an instability in the system. To our knowledge, a precise mechanism for such “hot electrons” has not been identified in the literature to date. Here we suggest another possible mechanism leading to the loss of a sharp insulating boundary. Namely, we propose that the electron layer loses stability due to a velocity shear type instability in crossed field devices, such as diocotron, magnetron, etc. In such an unstable layer, our observations show the formation of large-scale vortical structures that move in a stochastic way throughout the system (largely because of $\mathbf{E}\times\mathbf{B}$ drift), leading to stochastic oscillatory behavior in all quantities within the AK gap, including the conducting current. Such mechanism also results in non-zero loss current beyond the Hull field but does not necessarily produce a population of hot electrons. Ultimately, to be able to unambiguously identify the precise instability, a further study beyond the scope of this work is necessary.
@@
The stability properties of a parapotential magnetized electron layer have been discussed in the past [48, 49, 31]. While these properties depend on the layer’s spatial profile, of equal importance is whether and how this profile may in practice (experiment) be attained. Many factors make this problem very complex, however, for relevant to $Z$ parameters, even the simplest constant profile electron layer is never strictly stable, something recently emphasized in Ref. [39]. (See also reference [50], which considers vortex formation in geometries with an abrupt change in the AK gap size.) An attempt to create a stable electron layer by the authors of Ref. [51] was also not successful. Strong indication of a possible change in the stability properties of the electron layer in the AK gap can be seen in the analyses of the Hull curve [53, 52, 54]; for example, a jump in the SCL current around the value of near $B_{\rm crit}$ was noted in Ref. [53], while Refs. [52, 54] note the infinite slope of the curve at $B_{\rm crit}$. An earlier work by Pollack and Whinnery [55] observed both experimentally and computationally a sharp increase in the noise level in a planar diode as an externally imposed magnetic field exceeded the value of the Hull cutoff. We believe that in order to properly understand the loss of a sharp insulating boundary, a theory of current transport across an AK gap in unsteady flows must be developed.
@@
with positive $a$, $b$, and $\alpha$, and $X=B/B_{\rm crit}$ (in the Cartesian geometry, the magnetic field strength $B\equiv\left|B_{y}\right|$). Our Analytic fit Hull curve is obtained for values $a=0.065$, $b\in[1.03,1.04]$, and $\alpha=1$; the Circuit Element fit is obtained for $a=0.78$, $b=1.4$, and $\alpha=4$. It is seen that the function has a value of unity for zero magnetic field strength $B=0$. The notable feature here is that our analytic fit has a much smaller range outside of the critical magnetic field, exceeding it by only about 33334444%, compared to about $40$% for the Circuit Element fit.
Figure 8: Comparison of current losses in a 555\,5m long MITL with $I_{\rm peak}=20\,$MA/m, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm. The Circuit Element fit overpredicts EMPIRE 2D PIC result by about a factor of $3$ while the Analytic fit is in excellent agreement with it.
Since our PIC simulations were not successful in constructing the entire Hull curve, we have fitted the parameters in Eq. (21) from PIC simulations of current loss in MITLs of various geometrical dimensions. In particular, long MITLs have an amplifying effect on small variations in the curve parameter values. Fig. 8 shows a simulation of current loss in a 555\,5m long MITL ($I_{\rm peak}=20\,$MA/m, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm), which illustrates the difference in the predictions using the two different fits from Fig. 7. Clearly, the Circuit Element fit greatly overpredicts current losses, by a factor of about $3$. (The small difference between the 1D model using the Circuit Element fit and BERTHA is attributed to the fact that our model uses a fit, not the actual Hull curve used in BERTHA, see footnote 3.) Conversely, the details of the Hull curve in shorter MITLs in the length range $\lesssim 2\,$m do not manifest significant discrepancy with PIC simulations and both fits produce acceptable predictions.
One last remark in this section is that the “spillover” of a Hull curve acts as a low pass filter, and the larger the “spillover,” the stronger the low pass filter effect. For example, in Fig 8 we notice that the 1D model using the Analytic fit exhibits certain small scale temporal oscillations (see also Fig. 5). Such oscillations are not PIC noise since the 1D model is a (cold) fluid model. In the same figure, comparing with the simulations using the Circuit Element fit, we see a noticeably smoother curve. We recall that the Analytic fit exceeds $B_{\rm crit}$ by only about $3$-$4$% while the Circuit Element fit by about $40$%. This is not to imply that removing the small scale oscillations is nonphysical, rather, that it is possible that certain physical oscillations may unintentionally be filtered out.
@@ VI Temperature rise due to electron impact
We use the NIST stopping power tables for electrons in metals, $dK/dx$ [56]; or for certain metals, we prefer to use an analytic expression that gives the extra information in the low energy range [57]. First, the electron flux to the anode is directly related to the electron current:
@@
where $M_{\rm mol}$ is the molar mass of the metal, $\rho$ is the mass density, $C_{V}$ is the heat capacitance at constant volume, and $\theta_{i}$ is the electron angle of incidence measured from the normal to the surface (see also Ref. [18]). The total temperature increase is the sum over all time steps, $N_{\rm ins}=t_{\rm ins}/\Delta t$, until complete magnetic insulation is achieved,
@@
The only extra information necessary to calculate the temperature rise is the angle of incidence. It can be calculated as follows. Using the steady-state relations for conservation of the $z$-component of canonical momentum and the energy balance in Cartesian geometry [8, 58], we have
@@
Upon using Eq. (26), it is easy to recognize that setting the radical expression in Eq. (29) to zero, i.e., $\gamma^{2}-1-\left(eB_{y}d/mc\right)^{2}=0$, amounts to precisely the Hull condition [^6] The Hull condition can be derived from the same equations by setting $v_{x}=0$ at $x=d$, i.e., the condition that electrons do not reach the anode. (20). This means that the angle of incidence (29) is undefined at insulation (by Lovelace and Ott) as well as for larger values of the magnetic field, for which the expression under the square root becomes negative. In other words, using a Hull curve with a “spillover” is incompatible with the so outlined method of calculating the temperature. The simplest approach to resolving this incompatibility is to neglect temperature contributions at locations where $B_{y}\geq B_{\rm crit}$. This simple solution is acceptable for two reasons. First, we have seen from the Hull curve in Fig. 7 that the electron flux, which is proportional to the current, significantly drops near $B_{\rm crit}$, and so does its contribution to the temperature rise. And second, the neglected range of magnetic fields is only a small fraction of 33334444% of the total range for our Analytic fit. Our practical experience has confirmed that temperature doesn’t noticeably increase at locations where the magnetic field has exceeded the critical magnetic field.
@@
As a verification of this approach, Fig. 9 shows the temperature rise calculated from data from our 1D model and from CHICAGO 2D PIC, for aluminum metal initially at 300300300\,300K. The simulation setup is the same as in Fig. 4. To avoid small differences due to algorithmic implementation details, we use CHICAGO’s output quantities of current density and kinetic energy near the anode (quantities from particles information accumulated to the grid) to perform the calculations leading to Eq. (24). We notice that the overall agreement between the two models is excellent. This implies also that not only do both components of the current density agree very well (we have not shown $z$-components from either model) but so do the kinetic energies. The slight disagreement in the range of 20202020303030\,30cm at $t=9\,$ns (top panel) is due to the slight temporal delay of the PIC pulse compared to the 1D model, see Fig. 4 (bottom panel); however, just 111\,1ns later, at 101010\,10ns (bottom panel) no such disagreement is seen (the current loss pulses have caught up to one another). Since the temperature calculation is cumulative, such temporal offsets would not affect the final temperature increase.
@@
The obvious question concerns the large temperature oscillations in the PIC simulation in the first $\sim 10\,$cm of the MITL. These oscillations correspond to the current (and flux) oscillations in Fig. 4, which we believe to be a numerical artifact of the SCL emission algorithm in PIC, as discussed in Ref. [39].
@@
Figure 9: Comparison of temperature increase at $t=9\,$ns (top) and 101010\,10ns (bottom) between the 1D model and CHICAGO 2D PIC simulations, for $I_{\rm peak}=20\,$MA/m, $\tau_{\rm peak}=100\,$ns, AK gap, $d=5\,$mm, and $L=1\,$m.
@@ VII Conclusions
In this work we have discussed MITL losses in a strongly time-dependent setting. The physics that helps dramatically reduce current losses in long MITLs is that of the diamagnetic, non-local, cumulative electromagnetic effect of SCL currents. In short MITLs these effects diminish and in the limit vanish, and current losses can be predicted by the Child-Langmuir SCL law in the external (vacuum) EM fields. By comparing quantities of merit in systems with strong time dependence vs. those in (quasi-) steady-state, such as charge and current densities, current losses, etc., a more accurate definition of a long MITL is proposed.
A new 1D electromagnetic model is developed that performs on par with circuit element codes, with six to seven orders of magnitude lower computational cost than 2D PIC computations. The model is verified against 2D PIC simulations in a parallel plate Cartesian geometry as well as in a straight coaxial MITL geometry with azimuthal symmetry. This allows for efficient navigation of the large MITL design parameter space. A number of scaling studies have been performed, helping to understand and minimize MITL losses. A temperature diagnostic within the 1D model is easy to implement, based on the available information from the fields and currents. Agreement with the temperature calculations from PIC data further verifies that the 1D model accurately predicts kinetic energies and electron incidence angles as well.
A revised magnetic insulation model is proposed in the form a Hull curve, partly informed by direct PIC simulations and partly by fitting PIC simulation data. Major circuit element codes using a previously computed Hull curve tend to overpredict current losses in long MITLs but implementation of our revised Hull curve should not present difficulty.
It is proposed that previous experimental observations by Orzechowski and Bekefi of non-zero loss current beyond the theoretical critical magnetic field by Lovelace and Ott is due to the electron layer losing stability, not due to the generation of hot electrons, a physics picture consistent with our PIC simulations. This hypothesis still needs experimental confirmation.
Several extensions to our 1D model could be explored in the future. Including the transient motion of electrons across the AK gap could be achieved following calculations similar to those in Ref. [59]. Use of the relativistic Child-Langmuir law by Jory and Trivelpiece [60] should be straightforward, although the connection (16) would need revision. The inclusion of bipolar flow is another important extension left for the future. Thermally desorbed ions are frequently encountered in radially converging geometries and in many situations their contribution to current losses cannot be neglected. Thus, another extension to our model is to non-trivial geometries. Such opportunity is presented within our model by using geometrical transformations of fields and sources, for which well established techniques exist. Important geometric effects not captured by our 1D model — as well as by CEM models — are field enhancement around corners and the more general transverse magnetic (TM) modes in MITLs.